OpenStudy (anonymous):

Solve each equation for \[0\le x <2\pi\]. tan ^{2}x-(sqrt3+1)tanx+sqrt3 =0

OpenStudy (anonymous):

\[\tan ^{2}x-(\sqrt3+1)tanx+\sqrt3 =0\]

OpenStudy (anonymous):

Can you factor this?

OpenStudy (anonymous):

I don't know how to factor for this question. :(

OpenStudy (anonymous):

to*

OpenStudy (anonymous):

see tan(x) --> a for now

OpenStudy (anonymous):

What multiplies to make sqrt (3)? -sqrt(3) and -1? (tan x - ?)(tan x - ?)

OpenStudy (anonymous):

are you sure your expression is right?

OpenStudy (anonymous):

I think it is ...

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

can you guys show me step by step please?

OpenStudy (anonymous):

What about \[(\tan x - 1)(\tan x - \sqrt{3})\]

OpenStudy (anonymous):

that is right. But the question is how did you get there?

OpenStudy (anonymous):

a^2-(sqrt(3)+1)a+Sqrt[3] factored out to be (a -sqrt[3])(a -1)

OpenStudy (anonymous):

a= sqrt[3] or a= 1

OpenStudy (anonymous):

remember a=tan(x)

OpenStudy (anonymous):

I used the trial and error method of factoring from algebra tan x times tan x = tan^x -sqrt(3) times -1 = sqrt(3)

OpenStudy (anonymous):

tan[x]=sin(x)/cos(x)=sqrt[3]= sin(60)= sqrt[3]/2 cos(60)= 1/2 sqrt[3] --------- / ( 1/2) = sqrt[3] 2 hence x=60 degree

OpenStudy (anonymous):

I am only having trouble with the factoring in this question. I still don't get how you guys came up with (tanx-1)(tanx-sqrt3). I am sorry :(

OpenStudy (anonymous):

tan ^{2}x-(sqrt3+1)tanx+sqrt3 for simplicity, I substitute a=tan(x) a^2-(sqrt[3]+1)a+sqrt[3] then just factor

OpenStudy (anonymous):

Let me help with the factoring ... How would you factor x^2 - 5x + 6

OpenStudy (anonymous):

6 = -3 * -2 which means ( x - 3)(x - 2)??

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

oddly though i don't know how to factor in this question..

OpenStudy (anonymous):

If you know how to factor a^2 - 5a + 6 = (a - 3)(a-2) then factor tan ^{2}x-(sqrt3+1)tanx+sqrt3 =0 can we place the tan x where a is??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

And if -3 and -2 = 6 and -sqrt(3) and -1 = sqrt(3) then replace -3 and -2 with -sqrt(3) and -1

OpenStudy (anonymous):

(x - 3)(x -2) becomes (tan x - 1)(tan x - sqrt(3))

OpenStudy (anonymous):

success?

OpenStudy (anonymous):

you are wonderful! haha but I am still struggling.. mamammia.. :S

OpenStudy (anonymous):

one step at a time, you will make it .... the factor formed definitely looks weird and intimidating

OpenStudy (anonymous):

thanks! i am keep going over your steps and trying to make the way through! :)